### v1

# Exact treatment of the quantum Langevin equation under time-dependent system-bath coupling via a train of delta distributions

Abstract
In this paper, we consider the quantum Langevin equation for the Caldeira-Leggett model with an arbitrary time-dependent coupling constant. We solve this equation exactly by employing a train of Dirac-delta switchings. This method also enables us to visualize the memory effect in the environment. Furthermore, we compute the two-time correlation functions of the system’s quadratures and show that the discrete-time Fourier transform is well-suited for defining spectral densities, as the Dirac-delta switchings turn continuous functions into discretized samples.
I Introduction
The dissipative behavior of a quantum system interacting with its environment is one of the central features of open quantum systems [1, 2]. One of the key aspects in the study of open quantum systems is the memory effect in the environment. A Markovian dynamics is memoryless, which leads to one-way flow of information from a system to an environment. In contrast, a non-Markovian dynamics allows for the memory from the past interactions, which leads to a backflow of information from the environment to the system of interest [3].
The Caldeira-Leggett model [4, 5, 6] is a cornerstone in the theory of open quantum systems, describing the dynamics of a quantum Brownian particle [2]. In this framework, a quantum harmonic oscillator serves as the system, and it is coupled to an environment composed of a collection of quantum harmonic oscillators. Such a model is extremely useful for analyzing the behavior of open quantum systems because it allows for an exact solution to the equation of motion, known as the quantum Langevin equation (QLE) [1, 7, 8].
In many studies, the focus is on the asymptotic behavior of the system, e.g., the two-time correlation functions of the system quadratures in the long-interaction limit, during which the system reaches equilibrium with the environment. Therefore, it is assumed that the system constantly couples to the bath for an infinitely long time. In fact, this time-independent coupling is crucial for solving the QLE analytically. However, there have been few studies on the QLE with time-dependent coupling (i.e., non-stationary QLE). The primary reason is that obtaining an analytical solution is challenging. While some research applies Floquet theory to study the asymptotic behavior of periodically driven systems obeying the QLE [9, 10, 11, 12], the solution to the QLE with an arbitrary time-dependent coupling is still unknown. [^1] We note that, in the field of coarse-grained molecular dynamics simulation, the generalized Langevin equation (GLE) in the Mori-Zwanzig theory has been extensively studied. In recent years, non-stationary GLE has been analyzed by using the so-called memory reconstruction method [13, 14, 15]. This issue is particularly relevant in the study of finite-time interactions in, e.g., quantum thermodynamics [16].
To fill this gap, we employ a train of Dirac-delta switchings (a sum of Dirac’s delta distributions) to mimic a continuous time-dependent coupling, allowing us to solve the QLE analytically. This method is inspired by Refs. [17, 18], where a train of Delta switchings is employed to nonperturbatively analyze the behavior of qubits coupled to a quantum scalar field in curved spacetime. The Dirac delta distributions enable us to manage the dissipation kernel in the QLE, which is a major source of difficulty in solving it. Although one might initially think that a time-local delta distribution only produces the Markovian dynamics, we show that a collection of delta distributions can remarkably capture the non-Markovian behavior of the environment. In fact, the train of Dirac delta switchings enables us to pictorially understand how the memory effect in the environment plays a role. We show how to construct the diagram depicting the environment’s memory effect.
We also calculate the two-time correlation functions of the system’s quadratures, as well as the covariance matrix of the system. For the continuous constant coupling, one typically chooses a specific spectral density, such as the Lorentz-Drude spectral density, to evaluate correlation functions. These spectral densities are defined via the (continuous-time) Fourier transform of the dissipation kernel. For our delta switchings, however, the correlation functions are characterized by functions evaluated at each time where a delta-switching occurred. In other words, one deals with sequences instead of continuous functions to compute the correlation functions. Thus, the discrete-time Fourier transform is well-suited for defining spectral densities when a train of delta-switchings is adopted. We demonstrate that the correlation functions in the Dirac delta scenario converge to those for continuous switchings in the continuum limit.
This paper is organized as follows. In Sec. II, we review the QLE in the Caldeira-Leggett model. In Sec. III.1, we consider the QLE with a time-dependent coupling and solve it exactly by introducing the train of delta-switchings. The pictorial interpretation is given in III.2. We then introduce the concept of spectral density and study the two-time correlation functions in IV. The conventional spectral density is described in IV.1.1, and our spectral density defined via the discrete-time Fourier transform is introduced in IV.1.2. By using the spectral density for discrete-time samples, we demonstrate our result by computing the two-time correlation functions in IV.2. Throughout this paper, we set $\hbar=c=k_{\text{B}}=1$ and use the convention where the system’s mass is $M=1$.
II QLE with time-dependent interaction
In this section, we introduce the QLE with a time-dependent coupling. We begin by considering the Caldeira-Leggett model, where the environment is modeled as a collection of $\mathcal{N}$ independent quantum harmonic oscillators. Note that, at this stage, the environment contains a finite number of oscillators. After solving the QLE, we calculate the correlation functions and the covariance matrix of the system and then take the limit $\mathcal{N}\to\infty$.
Consider a system modeled by a quantum harmonic oscillator of frequency $\Omega$ (and mass $M=1$), whose quadratures are denoted by $Q$ and $P$. The free Hamiltonian $H_{\text{s,0}}$ is given by
$H_{\text{s,0}}$ $=\dfrac{P^{2}}{2}+\dfrac{\Omega^{2}}{2}Q^{2}\,.$ (1)
The environment is composed of $\mathcal{N}$ quantum harmonic oscillators, each labeled by $j\in\{1,2,\ldots,\mathcal{N}\}$. For each unit mass oscillator, the frequency and quadratures are given by $\omega_{j}$, $q_{j}$, and $p_{j}$, respectively, and the free Hamiltonian $H_{\text{E,0}}$ is
$H_{\text{E,0}}$ $=\sum_{j=1}^{\mathcal{N}}\left(\dfrac{p_{j}^{2}}{2}+\dfrac{\omega_{j}^{2}}{2}q_{j}^{2}\right).$ (2)
Assuming that the system couples identically to each oscillator in the environment, the interaction between the system and the environment is described by the interaction Hamiltonian
$H_{\text{int}}(t)$ $=-c(t)Q\otimes\sum_{j=1}^{\mathcal{N}}q_{j}\,,$ (3)
where $c(t)$ represents the time-dependent coupling between the system and the environment. For convenience, we write $c(t)=c\chi(t)$, where $c>0$ is the coupling constant and $\chi(t)$ is the switching function, which describes the time dependence of the interaction. In what follows, we assume that the interaction begins at time $t=0$.
To derive the QLE, consider the Heisenberg equations of motion for the quadratures in the Heisenberg picture. These lead to coupled second-order differential equations:
$\ddot{Q}(t)+\Omega^{2}Q(t)=c(t)\sum_{j=1}^{\mathcal{N}}q_{j}(t)\,,$ (4a)
$\ddot{q}_{j}(t)+\omega_{j}^{2}q_{j}(t)=c(t)Q(t)\,.$ (4b)
The equation of motion for oscillator-$j$ in (4b) can be solved as
$q_{j}(t)$ $=q_{j}^{(\text{h})}(t)$
$+\int_{0}^{\infty}\text{d}t^{\prime}\,\dfrac{\sin(\omega_{j}(t-t^{\prime}))}{\omega_{j}}\Theta(t-t^{\prime})c(t^{\prime})Q(t^{\prime})\,,$ (5)
where $q_{j}^{(\text{h})}(t)$ is the homogeneous solution to the equation, and $\Theta(t)$ is Heaviside’s step function. Inserting this solution into the right-hand side of (4a) gives the QLE:
$\ddot{Q}(t)+\Omega^{2}Q(t)+\int_{0}^{t}\text{d}t^{\prime}\,\Sigma(t,t^{\prime})Q(t^{\prime})$ $=\zeta(t)\,,$ (6)
where $\Sigma(t)$ is the so-called dissipation kernel (also known as the memory kernel) defined by
$\Sigma(t,t^{\prime})\coloneqq-\chi(t)\chi(t^{\prime})\Gamma(t-t^{\prime})\,,$ (7a)
$\Gamma(t)\coloneqq\sum_{j=1}^{\mathcal{N}}\dfrac{c^{2}}{\omega_{j}}\sin(\omega_{j}t)\,,$ (7b)
and the environment’s operator $\zeta(t)$, known as the noise term, is given by
$\zeta(t)$ $\coloneqq\chi(t)\xi(t)\,,\quad\xi(t)\equiv c\sum_{j=1}^{\mathcal{N}}q_{j}^{(\text{h})}(t)\,.$ (8)
We now comment on our derived QLE. Observe that, in general, the dissipation kernel $\Sigma(t,t^{\prime})$ is not time-translation invariant (i.e., nonstationary). An exception occurs when the switching function $\chi(t)$ is taken as a constant, $\chi(t)=1$, leading to $\Sigma(t,t^{\prime})\equiv\Sigma(t-t^{\prime})=-\Gamma(t-t^{\prime})$, so that the dissipation kernel only depends on the time difference (i.e., stationary). This particular case is widely employed in the literature, and the stationary QLE can be straightforwardly solved using the Laplace transformation. It is important to note that if the dissipation kernel is stationary, $\Sigma(t,t^{\prime})=\Sigma(t-t^{\prime})$, then the Laplace transform of the convolution integral is simply $\tilde{\Sigma}(z)\tilde{Q}(z)$ due to the convolution property. Here, $z\in\mathbb{C}$ is the Laplace variable, and $\tilde{Q}(z)\equiv\mathcal{L}[Q(t)]$, where $\mathcal{L}$ represents the Laplace transform with respect to $t$. In this case, the resulting algebraic equation in the Laplace domain leads to
$\tilde{Q}(z)$ $=z\tilde{G}(z)Q(0)+\tilde{G}(z)\dot{Q}(0)+\tilde{G}(z)\tilde{\zeta}(z)\,,$
where $\tilde{G}(z)\coloneqq(z^{2}+\Omega^{2}+\tilde{\Sigma}(z))^{-1}$. Notice that $\zeta(t)=\xi(t)$. The inverse Laplace transformation gives us the general solution
$Q(t)$ $=\dot{G}(t)Q(0)+G(t)\dot{Q}(0)+\int_{0}^{t}\text{d}t^{\prime}\,G(t-t^{\prime})\zeta(t^{\prime})\,.$ (9)
In practice, to compute quantities such as the correlation functions $\braket{Q(t)Q(t^{\prime})}$, one needs the explicit form of $G(t)$, the inverse Laplace transform of $\tilde{G}(z)$. This requires further assumptions, for example, specifying an explicit form of the spectral density (see, e.g., [19]), which we introduce later.
Nevertheless, if we do not assume a constant switching function, the dissipation kernel loses time-translation invariance, causing difficulties as we can no longer apply the convolution property. This is the main obstacle when considering an arbitrary switching function $\chi(t)$. In the following, we circumvent this issue by employing a collection of Dirac delta distributions.
III QLE with a train of Dirac delta
III.1 Solving the QLE
Suppose the switching function is compactly supported on $t\in[0,T]$. We choose the switching function in such a way that it is represented by $N$-Dirac delta distributions:
$\chi(t)$ $=\dfrac{T}{N}\sum_{k=1}^{N}\chi\left(\dfrac{k}{N}T\right)\delta\left(t-\dfrac{k}{N}T\right)$
$\equiv\dfrac{T}{N}\sum_{k=1}^{N}\chi(t_{k})\delta(t-t_{k})\,,$ (10)
where $t_{k}\equiv kT/N$, and the Dirac deltas are uniformly distributed between $t\in[0,T]$. Note that $T/N$ is the interval between successive interactions, $t_{k}$ and $t_{k+1}$. This allows us to simplify the term with the dissipation kernel in the QLE (6). Specifically, the Laplace transform of this term reads
$\mathcal{L}\int_{0}^{t}\text{d}t^{\prime}\,\Sigma(t,t^{\prime})Q(t^{\prime})$
$=-\dfrac{T^{2}}{N^{2}}\sum_{k,l=1}^{N}\chi(t_{k})\chi(t_{l})\Gamma(t_{k}-t_{l})\Theta(t_{k}-t_{l})Q(t_{l})e^{-zt_{k}}\,.$ (11)
The noise term $\chi(t)\xi(t)$ can also be transformed as
$\mathcal{L}[\zeta(t)]$ $=\sum_{k=1}^{N}\zeta(t_{k})e^{-zt_{k}}\,,\quad\zeta(t_{k})\equiv\dfrac{T}{N}\chi(t_{k})\xi(t_{k})\,.$ (12)
Therefore, after applying the Laplace transformation to the QLE and obtaining $\tilde{Q}(z)$, we perform the inverse Laplace transformation to obtain
$Q(t)$ $=Q_{t}^{(0)}+\sum_{l=1}^{N}K_{t,t_{l}}Q_{l}+\Xi_{t}\,,$ (13)
where
$Q_{t}^{(0)}\coloneqq Q(0)\cos(\Omega t)+\dfrac{P(0)}{\Omega}\sin(\Omega t)\,,$ (14a)
$K_{t,t_{l}}\equiv\dfrac{T^{2}}{N^{2}}\sum_{k=1}^{N}\Sigma(t_{k},t_{l})\Theta(t_{k}-t_{l})$
×$\qquad\times\dfrac{\sin[\Omega(t-t_{k})]}{\Omega}\Theta(t-t_{k})\,,$ (14b)
$Q_{l}\equiv Q(t_{l})\,,$ (14c)
$\Xi_{t}\equiv\sum_{k=1}^{N}\zeta(t_{k})\dfrac{\sin[\Omega(t-t_{k})]}{\Omega}\Theta(t-t_{k})\,.$ (14d)
Here, $Q_{t}^{(0)}$ is the solution to the Heisenberg equation of motion for a free quantum harmonic oscillator, $K_{t,t_{l}}$ is the term that emerges from the dissipation kernel (thus, responsible for the memory effect) with the property $K_{t,t_{N}}=0$ due to $\Sigma(t_{k},t_{k})=0$, $\Xi_{t}$ corresponds to the noise term, and $Q_{l}$ is $Q(t)$ at time $t=t_{l}$. Although $Q(t)$ in (13) takes a very simple form, it still depends on itself in the past, $Q_{l}$. Below, we obtain the general solution $Q(t)$ that depends only on the initial values $Q(0)$ and $P(0)$.
To this end, we express $Q_{l}$ in terms of $Q(0)$ and $P(0)$. Substituting $t=t_{l}$ in (13) gives us
$Q_{l}$ $=Q_{l}^{(0)}+\sum_{i=1}^{N}K_{li}Q_{i}+\Xi_{l}\,,$ (15)
where $Q_{l}^{(0)}$ and $\Xi_{l}$ are understood as $Q_{l}^{(0)}\equiv Q^{(0)}(t_{l})$ and $\Xi_{l}\equiv\Xi(t_{l})$. The equation above can be written in terms of vectors and matrices as follows:
$\bm{Q}$ $=\bm{Q}^{(0)}+\mathbf{K}\bm{Q}+\bm{\Xi}\,,$ (16)
where
$\bm{Q}$ $\coloneqq[Q_{1},Q_{2},\ldots,Q_{N}]^{\intercal}\,,$ (17a)
$\bm{Q}^{(0)}$ $\coloneqq[Q^{(0)}_{1},Q^{(0)}_{2},\ldots,Q^{(0)}_{N}]^{\intercal}\,,$ (17b)
$\mathbf{K}$ $\coloneqq\begin{bmatrix}0&0&0&\ldots&\ldots&0\\ 0&0&0&\ldots&\ldots&0\\ K_{31}&0&0&\ldots&\ldots&0\\ K_{41}&K_{42}&\ddots&\ddots&&\vdots\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ K_{N1}&K_{N2}&\ldots&K_{N,N-2}&0&0\\ \end{bmatrix}\,,$ (17c)
$\bm{\Xi}$ $\coloneqq[\Xi_{1},\Xi_{2},\ldots,\Xi_{N}]^{\intercal}\,.$ (17d)
Here, $K_{li}$ is the elements of the $N\times N$ matrix $\mathbf{K}$ given by
$K_{li}$ $=\dfrac{T^{2}}{N^{2}}\sum_{k=1}^{N}\Sigma(t_{k},t_{i})\Theta(t_{k}-t_{i})\dfrac{\sin[\Omega(t_{l}-t_{k})]}{\Omega}\Theta(t_{l}-t_{k})\,,$ (18)
and the matrix $\mathbf{K}$ is a strictly lower triangular matrix due to Heaviside’s step functions in $K_{li}$. Note that $K_{i+1,i}=0$ since $\Sigma(t_{k},t_{k})=0$.
The equation (16) can be solved for $\bm{Q}$ as
$\bm{Q}$ $=(\mathbf{I}-\mathbf{K})^{-1}\bm{Q}^{(0)}+(\mathbf{I}-\mathbf{K})^{-1}\bm{\Xi}\,,$ (19)
where $\mathbf{I}$ is the $N\times N$ identity matrix. Furthermore, the $N\times N$ strictly lower triangular matrix $\mathbf{K}$ has the nilpotent property: $\mathbf{K}^{N}=\mathbf{O}$, where $\mathbf{O}$ is the zero matrix. This allows us to express the inverse matrix $(\mathbf{I}-\mathbf{K})^{-1}$ as [^2] Here, $\mathbf{K}^{0}\equiv\mathbf{I}$.
$(\mathbf{I}-\mathbf{K})^{-1}$ $=\sum_{n=0}^{N-1}\mathbf{K}^{n}\equiv\mathsf{K}\,,$ (20)
because
$(\mathbf{I}-\mathbf{K})(\mathbf{I}+\mathbf{K}+\mathbf{K}^{2}+\ldots+\mathbf{K}^{N-1})$
$=\mathbf{I}-\mathbf{K}^{N}=\mathbf{I}-\mathbf{O}=\mathbf{I}\,.$ (21)
We thus obtained a compact form of the vector of $Q_{l}$:
$\bm{Q}$ $=\mathsf{K}\bm{Q}^{(0)}+\mathsf{K}\bm{\Xi}\,.$ (22)
Substituting $Q_{l}$ into $Q(t)$ in (13), we finally reached our main result:
$Q(t)$ $=\left[\cos(\Omega t)+\sum_{l,i=1}^{N}K_{t,t_{l}}\mathsf{K}_{li}\cos(\Omega t_{i})\right]Q(0)+\left[\dfrac{\sin(\Omega t)}{\Omega}+\sum_{l,i=1}^{N}K_{t,t_{l}}\mathsf{K}_{li}\dfrac{\sin(\Omega t_{i})}{\Omega}\right]P(0)$
$\quad+\sum_{k=1}^{N}\left[\dfrac{\sin[\Omega(t-t_{k})]}{\Omega}\Theta(t-t_{k})+\sum_{l,i=1}^{N}K_{t,t_{l}}\mathsf{K}_{li}\dfrac{\sin[\Omega(t_{i}-t_{k})]}{\Omega}\Theta(t_{i}-t_{k})\right]\zeta(t_{k})\,,$
$\equiv G[f_{Q}](t)Q(0)+G[f_{P}](t)P(0)+\sum_{k=1}^{N}G[f_{\zeta}^{(k)}](t)\zeta(t_{k})\,,\quad\forall t>0\,,$ (23)
where $G[f]$ is a functional defined as
$G[f](t)\coloneqq f(t)+\sum_{l,i=1}^{N}K_{t,t_{l}}\mathsf{K}_{li}f(t_{i})\,,$ (24a)
$f_{Q}(t)\coloneqq\cos(\Omega t)\,,$ (24b)
$f_{P}(t)\coloneqq\dfrac{\sin(\Omega t)}{\Omega}\,,$ (24c)
$f_{\zeta}^{(k)}(t)\coloneqq\dfrac{\sin[\Omega(t-t_{k})]}{\Omega}\Theta(t-t_{k})\,.$ (24d)
The function $G[f](t)$ in (24) explicitly shows how the memory effects of the environment affect the system. If the dissipation kernel is negligible, then $K_{t,t_{l}}$ and $\mathsf{K}_{li}$ vanish. In what follows, we denote $G_{f}(t)\equiv G[f](t)$ for brevity.
Equation (23) is the exact solution to the QLE (6) when a compactly supported switching function is described by the train of Dirac deltas (10). Note that this solution is also valid for intermediate times $t\in(0,T)$, as the Heaviside step functions naturally eliminate irrelevant terms. Moreover, at each discrete time $t_{l}$, $l\in\{1,2,\ldots,N\}$, the solution $Q(t=t_{l})$ reduces to Eq. (22), which can be expressed using a single function $\mathcal{G}(t)$ as
$Q(t_{l})$ $=\dot{\mathcal{G}}(t_{l})Q(0)+\mathcal{G}(t_{l})P(0)$
$\quad+\sum_{k=1}^{N}\mathcal{G}(t_{l}-t_{k})\Theta(t_{l}-t_{k})\zeta(t_{k})\,,$ (25)
where
$\mathcal{G}(t)$ $\coloneqq\dfrac{\sin(\Omega t)}{\Omega}+\sum_{l,i=1}^{N}K_{t,t_{l}}\mathsf{K}_{li}\dfrac{\sin(\Omega t_{i})}{\Omega}\,.$ (26)
This is consistent with the well-known solution for constant switching, as the continuum limit of Eq. (25) reduces to Eq. (9).
Figure 1: The basic diagrams that constitute (27). Each $t_{i}$, $i\in\{1,2,\ldots,N\}$ represents the time at which the instantaneous interaction occurs. Here, we use a simplified notation $\Sigma_{ki}\equiv\Sigma(t_{k},t_{i})$.
III.2 Pictorial interpretation
Each term in our result (23) consists of time-dependent functions $G_{f_{Q}}(t),G_{f_{P}}(t)$, and $G_{f_{\zeta}}(t)$, which are given by (24). These functions are written in the form
$G_{f}(t)$ $=f_{t}+\sum_{l,i=1}^{N}K_{t,t_{l}}\mathsf{K}_{li}f_{i}\,,$ (27)
where $f_{t}\equiv f(t)$. This expression clearly shows the influence of the dissipation kernel, as both $K_{t,t_{l}}$ and $\mathsf{K}_{li}$ contain $\Sigma(t)$. Although the second term appears cumbersome, we show that it can be intuitively understood using diagrams.
Figure 2: The diagrams illustrating $G_{f}(t)$ in (27) for (a) $N=2$, (b) $N=3$, and (c) $N=4$. A loop connecting $t_{i}$ and $t_{j}$ ($j>i$) corresponds to $\Sigma(t_{j},t_{i})\equiv\Sigma_{ji}$.
To begin with, we introduce diagrams that characterize each term in (27). The function $f_{t}$ describes $G_{f}(t)$ in the absence of the dissipation kernel. We depict $f_{t}$ as a dashed line along the time axis, as shown in Fig. 1(i). In particular, $f_{i}\equiv f(t_{i})$ is illustrated as a dashed line from $t=0$ to $t=t_{i}$ [Fig. 1(ii)]. The quantity $K_{t,t_{l}}$, defined in (14b), is the sum of the products of $\Sigma(t_{k},t_{l})\Theta(t_{k}-t_{l})$ and $\sin[\Omega(t-t_{k})]\Theta(t-t_{k})/\Omega$. The factor $\Sigma(t_{k},t_{l})\Theta(t_{k}-t_{l})$ is interpreted as an “exchange” of a signal from $t=t_{l}$ to $t_{k}$ due to the environment’s memory effect. Subsequently, the term $\sin[\Omega(t-t_{k})]\Theta(t-t_{k})/\Omega$ propagates this effect from $t=t_{k}$ to $t(>T)$. Thus, we assign a solid loop to $\Sigma(t_{k},t_{l})$ and a line to $\sin[\Omega(t-t_{k})]\Theta(t-t_{k})/\Omega$ as illustrated in Fig. 1(iii). A similar interpretation applies to $K_{li}$ defined in (18), except that the propagation terminates at $t=t_{l}$ [Fig. 1(iv)].
We note that $\Sigma(t_{k},t_{k})=0$ for any environment. This means that an “instantaneous loop”, which is closed at a single time $t=t_{k}$, is not allowed. Physically, it guarantees that there is no immediate backreaction from the environment to the system.
Let us use these diagrams to understand the role of memory effects. We first note that the elements of the matrix $\mathsf{K}$ consist of $K_{li}$. For instance,
$\mathsf{K}$ $=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\quad\text{for }N=2\,,$
$\mathsf{K}$ $=\begin{bmatrix}1&0&0\\ 0&1&0\\ K_{31}&0&1\end{bmatrix}\quad\text{for }N=3\,,$
$\mathsf{K}$ $=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ K_{31}&0&1&0\\ K_{41}&K_{42}&0&1\end{bmatrix}\quad\text{for }N=4\,,$
$\mathsf{K}$ $=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ K_{31}&0&1&0&0\\ K_{41}&K_{42}&0&1&0\\ K_{51}+K_{53}K_{31}&K_{52}&K_{53}&0&1\end{bmatrix}\quad\text{for }N=5\,.$
Let us take a simple example of $N=2$ and visualize $G_{f}(t)$ in (27). Reminding that $K_{t,t_{N}}=0$, the function $G_{f}(t)$ for $N=2$ can be explicitly written as
$G_{f}(t)$ $=f_{t}+\sum_{l,i=1}^{N=2}K_{t,t_{l}}\mathsf{K}_{li}f_{i}$
$=f_{t}+\begin{bmatrix}K_{t,t_{1}}&0\end{bmatrix}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{bmatrix}f_{1}\\ f_{2}\end{bmatrix}$
$=f_{t}+K_{t,t_{1}}f_{1}\,.$ (28)
Figure 2(a) illustrates this expression using our diagrammatic elements. The term $K_{t,t_{1}}f_{1}$ can be understood as follows. First, draw a dashed line representing $f_{1}$ from $t=0$ to $t_{1}$. Then, we consider the factor $K_{t,t_{1}}$, which contains a loop and a line as illustrated in Fig. 1(iii). In general, all possible configurations must be considered for each term. In the $N=2$ case, there is only one configuration for $K_{t,t_{1}}f_{1}$ as shown in Fig. 2(a). For $N=3$, however, there are two possible configurations for $K_{t,t_{1}}f_{1}$ as depicted in Fig. 2(b), and their sum gives the final value of $K_{t,t_{1}}f_{1}$.
For $N=4$, various configurations emerge [Fig. 2(c)]. The same argument above applies to $K_{t,t_{2}}f_{2}$ and $K_{t,t_{3}}f_{3}$. We also have the term $K_{t,t_{3}}K_{31}f_{1}$, which contains two loops arising from $K_{t,t_{3}}$ and $K_{31}$. Here, $K_{31}$ is represented by a loop starting from $t_{1}$ and a line ending at $t_{3}$.
Overall, the expression for $G_{f}(t)$ sums over all possible memory effects of the environment. We stress that both Markovian and non-Markovian effects can also be understood using the diagrams. Markovian processes correspond to diagrams that consist of a single loop connecting adjacent times, namely, the terms with $\Sigma(t_{i+1},t_{i})$. One can be convinced by considering the “continuum limit” of the delta switchings, $N\to\infty$ and $t_{i+1}-t_{i}(\equiv T/N)\to 0$. In this case, the adjacent memory effect $\Sigma(t_{i+1},t_{i})$ can be considered time-local. On the other hand, non-Markovian effects are represented by loops connecting distant points $t_{i}$ and $t_{i+j}$ with $j\geq 2$. In Fig. 2(c), the Markovian diagrams correspond to $\Sigma_{21}$, $\Sigma_{32}$, and $\Sigma_{43}$ in $K_{t,t_{1}}f_{1}$, $K_{t,t_{2}}f_{2}$, and $K_{t,t_{3}}f_{3}$, respectively. Note that the diagram corresponding to $K_{t,t_{3}}K_{31}f_{1}$, which contains two adjacent loops, is not interpreted as Markovian. The mathematical argument is provided in Appendix B.
III.3 Two-time correlation functions and the covariance matrix
In the literature, two-time correlation functions such as $\braket{Q(t)Q(t^{\prime})}$ and $\braket{P(t)P(t^{\prime})}$, as well as their asymptotic values are the primary focus. A related quantity of interest, especially in Gaussian quantum mechanics [20], is the covariance matrix $\mathbb{V}$, which determines a quantum state. In the following, we study these by using our solution to the QLE (23).
III.3.1 Two-time correlation functions
In the Heisenberg picture, consider the system’s quadrature $\bm{R}(t)\equiv[Q(t),P(t)]^{\intercal}$. The two-time correlation function is defined by
$\braket{R_{i}(t)R_{j}(t^{\prime})}$ $\equiv\braket{R_{i}(t)R_{j}(t^{\prime})}_{\rho_{\text{tot}}(0)}$
$\coloneqq\operatorname{Tr}[\rho_{\text{tot}}(0)R_{i}(t)R_{j}(t^{\prime})]\quad(i,j\in\{1,2\})\,.$ (29)
We note that the expectation values for $\bm{R}(t)$ are taken with respect to the total initial state $\rho_{\text{tot}}(0)$, as $Q(t)$ and $P(t)$ are considered to be observables on the total Hilbert space [see Eq. (23)]. On the other hand, the expectation values for the initial quadratures $Q(0)$ and $P(0)$, and those for the environment’s observable $\zeta(t)\equiv\chi(t)\xi(t)$ should be understood as $\braket{R_{j}(0)}\equiv\braket{R_{j}(0)}_{\rho_{\text{s}}(0)}$ and $\braket{\zeta(t)}\equiv\braket{\zeta(t)}_{\rho_{\text{E}}(0)}$, where $\rho_{\text{s}}(0)$ and $\rho_{\text{E}}(0)$ are the initial states of the system and the environment, respectively. In what follows, we omit the subscripts $\rho_{\text{tot}}(0)$, $\rho_{\text{s}}(0)$, and $\rho_{\text{E}}(0)$.
Let us focus on $\braket{Q(t)Q(t^{\prime})}$ and explicitly write in terms of the functions in (24). We assume that the initial joint state is a product state,
$\rho_{\text{tot}}(0)=\rho_{\text{s}}(0)\otimes\rho_{\text{E}}(0)\,,$ (30)
and that the environment’s one-point correlation function is zero, $\braket{\zeta(t)}=0$, and any $n$-point correlation functions, $\braket{\zeta(t_{1})\ldots\zeta(t_{n})}$, can be written as products of the environment’s two-time correlation functions, $\braket{\zeta(t)\zeta(t^{\prime})}$. The thermal Gibbs state is one of the examples. From (23), the two-time correlation function $\braket{Q(t)Q(t^{\prime})}$ reads
$\braket{Q(t)Q(t^{\prime})}$ $=G_{f_{Q}}(t)G_{f_{Q}}(t^{\prime})\braket{Q^{2}(0)}$
$\quad+G_{f_{P}}(t)G_{f_{P}}(t^{\prime})\braket{P^{2}(0)}$
$\quad+G_{f_{Q}}(t)G_{f_{P}}(t^{\prime})\braket{Q(0)P(0)}$
$\quad+G_{f_{P}}(t)G_{f_{Q}}(t^{\prime})\braket{P(0)Q(0)}$
$\quad+\sum_{k,k^{\prime}=1}^{N}G_{f_{\zeta}^{(k)}}(t)G_{f_{\zeta}^{(k^{\prime})}}(t^{\prime})\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}\,.$ (31)
As we point out later, the delta switchings allow us to write the correlation functions in terms of discretized data, $\Gamma(t_{k})$ and $\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}$. This suggests that the discrete-time Fourier transform is preferred (as opposed to the continuous-time Fourier transform) for defining quantities in the frequency domain such as the spectral density.
III.3.2 The covariance matrix
In continuous-variable quantum mechanics, a quantum state is called Gaussian if it is determined solely by the first and second moments. We denote the first moment by $\bar{\bm{R}}\equiv\braket{\bm{R}}$. For the second moments, one typically considers the covariance matrix—a real symmetric, positive-definite matrix—$\mathbb{V}(t)$ for describing the second moments, and it is given by [21]
$\mathbb{V}_{ij}(t)$$\coloneqq\braket{\{R_{i}(t),R_{j}(t)\}}-2\braket{R_{i}(t)}\braket{R_{j}(t)}\,.$ (32)
The vacuum state, the thermal Gibbs state, and the squeezed state are examples of the Gaussian state with vanishing first moments, and the coherent state is a Gaussian state with a nonvanishing first moment. Once these statistical moments are known, we can obtain the corresponding Gaussian state $\rho_{\text{G}}\equiv\rho_{\text{G}}[\bar{\bm{R}},\mathbb{V}]$.
A particularly useful tool for studying the covariance matrix in Gaussian systems is the Gaussian operations. These are the completely-positive (CP) maps that transform Gaussian states to Gaussian states [20]. To illustrate this, suppose a joint system is initially prepared in a product of Gaussian states, $\rho_{\text{tot}}(0)=\rho_{\text{s}}(0)\otimes\rho_{\text{E}}(0)$, which is also Gaussian. A unitary time-evolution operator generated by a Hamiltonian composed of a linear and a quadratic term in $R_{i}$ is an example of the Gaussian operation known as the Gaussian unitary transformation [20]. The Hamiltonian in the Caldeira-Leggett model used in this paper indeed generates such a unitary transform, and this fact does not change when a train of delta-switchings is employed. Moreover, a partial trace of a Gaussian joint state is also a Gaussian operation. Thus, the system’s reduced quantum state $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U(t)\rho_{\text{tot}}(0)U^{\dagger}(t)]$ after the time-evolution remains to be Gaussian. This fact can be utilized to characterize the evolution of the statistical moments. Assuming the initial joint state is separable Gaussian and they evolve under a trace-preserving Gaussian operation (hence, it is a CPTP map), the first moment and the covariance matrix evolve as [20]
$\bar{\bm{R}}\mapsto\mathbb{T}\bar{\bm{R}}\,,$ (33a)
$\mathbb{V}\mapsto\mathbb{T}\mathbb{V}\mathbb{T}^{\intercal}+\mathbb{N}\,,$ (33b)
where $\mathbb{T}$ and $\mathbb{N}$ are real square matrices satisfying the condition that emerges from the uncertainty principle,
$\mathbb{N}+\mathsf{i}\bm{\Omega}\geq\mathsf{i}\mathbb{T}\bm{\Omega}\mathbb{T}^{\intercal}\,,$ (34)
where $\bm{\Omega}$ is the symplectic form that appears in the canonical commutation relations: $[R_{i},R_{j}]=\mathsf{i}\Omega_{ij}$.
One can prove that the Markovianity of a quantum channel is encoded in the properties of $\mathbb{T}$ and $\mathbb{N}$. Let $\mathcal{E}_{t,s}:\rho(s)\mapsto\rho(t)$ be a Gaussian CPTP map. If $\mathcal{E}_{t,s}$ is a Markovian dynamical map, then the associated matrices $\mathbb{T}$ and $\mathbb{N}$ obey:
$\mathbb{T}_{t+s}$ $=\mathbb{T}_{t}\mathbb{T}_{s}\,,\quad\mathbb{N}_{t+s}=\mathbb{T}_{t}\mathbb{N}_{s}\mathbb{T}_{t}^{\intercal}+\mathbb{N}_{t}\,,$ (35)
where $\mathbb{T}_{t}$ and $\mathbb{N}_{t}$ are the matrices associated with $\mathcal{E}_{t,0}$. We prove these properties in Appendix A.
The evolution for $\bar{\bm{R}}$ and $\mathbb{V}$ given in Eq. (33) holds for generic CPTP Gaussian operations, as long as the initial joint state is separable Gaussian. We now apply this to our delta-switching case when the system is initially prepared in a Gaussian state. In particular, we assume that the environment’s initial Gaussian state has a zero first moment, $\braket{\zeta(0)}=0$.
From the result in (31) and other correlation functions such as $\braket{\{Q(t),P(t)\}}$ allow us to obtain the matrices $\mathbb{T}_{t}$ and $\mathbb{N}_{t}$ for $t>T$ as
$\mathbb{V}(t)=\mathbb{T}_{t}\mathbb{V}(0)\mathbb{T}^{\intercal}_{t}+\mathbb{N}_{t}\,,$
$\mathbb{T}_{t}=\begin{bmatrix}G_{f_{Q}}(t)&G_{f_{P}}(t)\\ \dot{G}_{f_{Q}}(t)&\dot{G}_{f_{P}}(t)\end{bmatrix}\,,$ (36)
$\mathbb{N}_{t}=\begin{bmatrix}N_{QQ}(t)&N_{QP}(t)\\ N_{PQ}(t)&N_{PP}(t)\end{bmatrix},$ (37)
where
$N_{QQ}(t)$ $=\sum_{k,k^{\prime}=1}^{N}G_{f_{\zeta}^{(k)}}(t)G_{f_{\zeta}^{(k^{\prime})}}(t)\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}\,,$ (38a)
$N_{QP}(t)$ $=\sum_{k,k^{\prime}=1}^{N}G_{f_{\zeta}^{(k)}}(t)\dot{G}_{f_{\zeta}^{(k^{\prime})}}(t)\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}\,,$ (38b)
$N_{PQ}(t)$ $=\sum_{k,k^{\prime}=1}^{N}\dot{G}_{f_{\zeta}^{(k)}}(t)G_{f_{\zeta}^{(k^{\prime})}}(t)\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}\,,$ (38c)
$N_{PP}(t)$ $=\sum_{k,k^{\prime}=1}^{N}\dot{G}_{f_{\zeta}^{(k)}}(t)\dot{G}_{f_{\zeta}^{(k^{\prime})}}(t)\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}\,.$ (38d)
It turns out that the matrices $\mathbb{T}_{t}$ and $\mathbb{N}_{t}$ for the case of delta-switchings do not satisfy the Markovianity conditions (35) in general. However, one can reduce the dynamics to Markovian by applying the Born-Markov approximation. We demonstrate this fact in Appendix B for an environment with the so-called Lorentz-Drude spectral density described in the next section.
IV Demonstration with the Lorentz-Drude spectral density
IV.1 Spectral density
Our main result in (23) was derived under the assumption that the environment consists of a finite collection of quantum harmonic oscillators. Nevertheless, we can extend this result to the continuum limit by introducing the spectral density $\sigma(\omega)$ of the environment. In the traditional QLE with a continuous constant coupling, the spectral density is defined via the Fourier transform of the dissipation kernel. In contrast, we illustrate that the spectral density should be defined by using the discrete-time Fourier transform (DTFT) when a train of delta switchings is employed.
IV.1.1 Introducing the spectral density
Let us introduce the traditional spectral density used when a continuous constant coupling is considered. Our first aim is to rewrite the quantities that appear in $\braket{R_{i}(t)}$ and $\braket{R_{i}(t)R_{j}(t^{\prime})}$ in terms of the spectral density $\sigma(\omega)$. Consider the time-translation invariant part $\Gamma(t)$ of the dissipation kernel. For a finite collection of quantum harmonic oscillators, $\Gamma(t)$ is given by (7b). We then introduce the spectral density $\sigma(\omega)$ by Fourier transforming $\Gamma(t)$ as [^3] In this paper, the Fourier transform is denoted by $\hat{\Gamma}(\omega)\equiv\mathcal{F}[\Gamma(t)]$. This notation should not be confused with that used for linear operators in quantum mechanics.
$\hat{\Gamma}(\omega)$ $\coloneqq\int_{\mathbb{R}}\text{d}t\,\Gamma(t)e^{\mathsf{i}\omega t}\equiv 2\mathsf{i}\sigma(\omega)\,,$ (39)
where
$\sigma(\omega)$ $=\sum_{j=1}^{\mathcal{N}}\dfrac{\pi}{2}\dfrac{c^{2}}{\omega_{j}}\left[\delta(\omega-\omega_{j})-\delta(\omega+\omega_{j})\right]$ (40)
is the spectral density of the environment composed of a finite collection of harmonic oscillators. Note that it has the property $\sigma(\omega)=-\sigma(-\omega)$. Then, $\Gamma(t)$ can be expressed using the spectral density as
$\Gamma(t)$ $=\dfrac{\mathsf{i}}{\pi}\int_{\mathbb{R}}\text{d}\omega\,\sigma(\omega)e^{-\mathsf{i}\omega t}\,,$ (41)
and thereby the dissipation kernel $\Sigma(t,t^{\prime})$ reads
$\Sigma(t,t^{\prime})$ $=-\chi(t)\chi(t^{\prime})\dfrac{\mathsf{i}}{\pi}\int_{\mathbb{R}}\text{d}\omega\,\sigma(\omega)e^{-\mathsf{i}\omega(t-t^{\prime})}\,.$ (42)
We also need to express the quantities related to the noise term $\zeta(t)$, defined in (8), in terms of $\sigma(\omega)$. Instead of establishing a direct relationship between $\zeta(t)$ and $\sigma(\omega)$, we express the correlation functions $\braket{\zeta(t)\zeta(t^{\prime})}$ in terms of $\sigma(\omega)$, as these correlations are required for evaluating the two-time correlation functions. To this end, we assume that the initial joint state is a product state and that the environment is prepared in a thermal Gibbs state:
$\rho_{\text{tot}}(0)$ $=\rho_{\text{s}}(0)\otimes\rho_{\text{E}}(0)\,,\quad\rho_{\text{E}}(0)=\dfrac{1}{Z}e^{-\beta H_{\text{E}}}\,,$ (43)
where $Z\coloneqq\operatorname{Tr}[e^{-\beta H_{\text{E}}}]$ is the partition function and $\beta>0$ is the inverse temperature of the environment. Then, one can straightforwardly verify that the one-point and two-point correlation functions read
$\braket{\zeta(t)}=0\,,$ (44a)
$\braket{\zeta(t)\zeta(t^{\prime})}=\chi(t)\chi(t^{\prime})\braket{\xi(t)\xi(t^{\prime})}\,,$ (44b)
$\braket{\xi(t)\xi(t^{\prime})}=\dfrac{1}{\pi}\int_{\mathbb{R}}\text{d}\omega\,\dfrac{\sigma(\omega)}{e^{\beta\omega}-1}e^{-\mathsf{i}\omega(t-t^{\prime})}\,.$ (44c)
Again, the correlation functions for the environment’s observable should be understood as $\braket{\zeta(t)}\equiv\braket{\zeta(t)}_{\rho_{\text{E}}(0)}$.
By introducing the spectral density $\sigma(\omega)$, each term in the two-time correlation function can be rewritten in terms of it. This approach allows us to examine a variety of environments, including those with an infinite number of quantum harmonic oscillators. The idea is as follows. So far, our $\sigma(\omega)$ given in (40) is the explicit form for a finite collection of quantum harmonic oscillators. To consider other types of environments (e.g., the case where $\mathcal{N}\to\infty$), we simply replace the spectral density in (40) with the one of our interest. A widely examined spectral density is the Lorentz-Drude (LD) spectral density:
$\sigma(\omega)$ $=\gamma\dfrac{\omega\Lambda^{2}}{\omega^{2}+\Lambda^{2}}\,,$ (45)
where $\gamma>0$ is the coupling constant and $\Lambda>0$ is the cutoff frequency that determines the bandwidth of the environment. Moreover, $\Lambda^{-1}$ is the characteristic time scale of the change in the environment. Given a cutoff $\Lambda$, the low-frequency regime $\omega\ll\Lambda$ of the LD spectral density mimics the Ohmic system as $\sigma(\omega)\approx\gamma\omega$, which reflects the Markovian dynamics. Employing the LD spectral density, we have
$\Gamma(t)$ $=\gamma\Lambda^{2}e^{-\Lambda|t|}\text{sgn}(t)\,.$ (46)
IV.1.2 Spectral density for discretized data
When considering continuous (typically constant) switching functions $\chi(t)$, one can compute correlation functions using the spectral density defined above. In this paper, however, the train of delta switchings requires us to adopt the discrete-time Fourier transform (DTFT) [22]—the discrete-time variant of the continuous-time Fourier transform—to define the spectral density.
To see this, consider the two-time correlation functions such as $\braket{Q(t)Q(t^{\prime})}$. As we saw in (31), these correlation functions are composed of the terms such as $\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}$, which is the noise correlation function evaluated at the delta-switched (discrete) times $t_{k}$ and $t_{k^{\prime}}$. Therefore, instead of employing the continuous-time Fourier transform to define the spectral density and the noise correlations, it is suitable to choose the DTFT for our delta-switched scenario. [^4] Simply inserting $t=t_{k}$ and $t^{\prime}=t_{k^{\prime}}$ in the noise correlator (44c) defined using the continuous-time Fourier transform leads to divergence at $k=k^{\prime}$.
From now on, we redefine the spectral density $\sigma(\omega)$ by employing the DTFT. Let $\{f_{n}\}$ be an absolutely summable discrete-time sequence. The DTFT of $f_{k}$ [denoted by $\breve{f}(\overline{\omega})$] is defined as
$\breve{f}(\overline{\omega})$ $\coloneqq\sum_{n=-\infty}^{\infty}f_{n}e^{\mathsf{i}\overline{\omega}n}\,,$ (47)
and the inverse DTFT is given by
$f_{n}$ $=\int_{-\pi}^{\pi}\dfrac{\text{d}\overline{\omega}}{2\pi}\,\breve{f}(\overline{\omega})e^{-\mathsf{i}\overline{\omega}n}\,.$ (48)
Note that $\overline{\omega}$ has units of radians, which relates to the frequency $\omega$ in the Fourier transform via $\overline{\omega}=\omega T/N$. Also $\breve{f}(\overline{\omega})$ is periodic, $\breve{f}(\overline{\omega}+2\pi)=\breve{f}(\overline{\omega})$, where the periodicity $2\pi$ is called the sampling frequency. Recalling that a periodic function can be expressed as a Fourier series, the definition of the DTFT (47) is basically the Fourier series in the frequency domain, and $f_{n}$ corresponds to the Fourier coefficient.
Using the DTFT defined above, let us consider $\Gamma_{k}\equiv\Gamma(t_{k})$. Our aim is to replace the spectral density $\sigma(\omega)$ with the periodic spectral density $s(\overline{\omega})$ associated with $\Gamma_{k}$ as shown in Eq. (53). The idea is as follows. Suppose the environment of interest has the memory effect described by the LD spectral density (45), which gives $\Gamma(t)$ in (46). The train of delta-switchings allows us to compute the two-time correlation functions by using discretized data such as $\Gamma_{k}$ and
$\braket{\zeta_{k}\zeta_{k^{\prime}}}\equiv\braket{\zeta({t_{k}})\zeta(t_{k^{\prime}})}=\dfrac{T^{2}}{N^{2}}\chi(t_{k})\chi(t_{k^{\prime}})\braket{\xi_{k}\xi_{k^{\prime}}}\,.$ (49)
Here $\braket{\xi_{k}\xi_{k^{\prime}}}\equiv\braket{\xi({t_{k}})\xi(t_{k^{\prime}})}$ and we used Eq. (12). Thus, it is natural to employ the DTFT to express the frequency representations of these discretized data, and we denote the spectral density defined through the DTFT by $s(\overline{\omega})$.
Figure 3: Comparison between the solutions to the QLE with a constant switching (dashed line) and our train of delta switchings (solid curve). The horizontal axis is taken to be the number of delta switchings $N$. Here, we choose $\gamma/\Omega=0.1,\Lambda/\Omega=2$. (a) The function $G_{f_{P}}$ at fixed time $T=1$ asymptotes to $G(t=1)$ as the number of delta switchings $N$ increases. (b) The correlation function evaluated with the train of delta switchings also asymptotes to the result for constant switching.
For concreteness, let us insert $t=t_{k}$ into $\Gamma(t)$ in (46):
$\Gamma_{k}$ $=\gamma\Lambda^{2}e^{-\Lambda T|k|/N}\text{sgn}(k)\,.$ (50)
The DTFT of $\Gamma_{k}$ reads [^5] Recall that we initially introduced a train of $N$ delta switchings from $k=1$ to $k=N$ [see (10)]. In order to employ the DTFT, we extended $k$ from $k\in\{1,2,\ldots,N\}$ to $k\in\mathbb{Z}$ by assuming that the amplitude of switching $\chi(t_{k})$ is zero outside the support $t\in[0,T]$.
$\breve{\Gamma}(\overline{\omega})$ $=\sum_{k=-\infty}^{\infty}\Gamma_{k}e^{\mathsf{i}\overline{\omega}k}=\dfrac{-\mathsf{i}\gamma\Lambda^{2}\sin\overline{\omega}}{\cos\overline{\omega}-\cosh(\Lambda T/N)}\,,$ (51)
and we define our spectral density $s(\overline{\omega})$ for the discretized data by $2\mathsf{i}s(\overline{\omega})\equiv\breve{\Gamma}(\overline{\omega})$ analogous to (39) so that
$s(\overline{\omega})$ $=-\dfrac{1}{2}\dfrac{\gamma\Lambda^{2}\sin\overline{\omega}}{\cos\overline{\omega}-\cosh(\Lambda T/N)}\,.$ (52)
We again stress that $s(\overline{\omega})$ is $2\pi$-periodic.
The inverse DTFT gives us the relation
$\Gamma_{k}$ $=\dfrac{\mathsf{i}}{\pi}\int_{-\pi}^{\pi}\text{d}\overline{\omega}\,s(\overline{\omega})e^{-\mathsf{i}\overline{\omega}k}\,.$ (53)
Applying the same logic used in (44c), the discretized noise correlation function in the thermal state is given by
$\braket{\xi_{k}\xi_{k^{\prime}}}$ $=\dfrac{1}{\pi}\int_{-\pi}^{\pi}\text{d}\overline{\omega}\,\dfrac{s(\overline{\omega})}{e^{\beta\overline{\omega}}-1}e^{-\mathsf{i}\overline{\omega}(k-k^{\prime})}\,.$ (54)
Summarizing, we use the LD spectral density $\sigma(\omega)$ in (45) and the associated quantities like (44c) and (46) when the switching function is continuous. On the other hand, when we employ the train of delta-switchings, we instead use the $2\pi$-periodic spectral density $s(\overline{\omega})$ in (52) and the related discretized data (54). These two spectral densities give essentially the same memory effects in the time-domain, (46) and (50), respectively, and they are related by the Poisson summation formula:
$s(\omega T/N)$ $=\dfrac{N}{T}\sum_{k=-\infty}^{\infty}\sigma(\omega-2\pi kN/T)\,.$ (55)
IV.2 Comparison to constant switching
We numerically demonstrate that our solution (23) to the QLE using the delta switchings agrees with the well-known result of the continuous constant switching (9).
Consider the well-known solution (9) to the QLE with $\chi(t)=1$. We choose the LD spectral density $\sigma(\omega)$ given in Eq. (45) and evaluate $G(t)$ and $\braket{Q^{2}(t)}$. We then compare these to our solution (23) with the DTFT-version of LD spectral density $s(\omega)$ given in Eq. (52). Here, we choose $\chi(t_{k})=1$ for all $k\in\{1,2,\ldots,N\}$.
Figure 3(a) shows $G_{f_{P}}(T)$ evaluated against the number of delta switchings $N$. Here, we choose $T=1$. As the number of switching times $N$ increases, our $G_{f_{P}}(T)$ asymptotes to the well-known continuous solution $G(t)$ at $t=1$. This indicates that the delta-switching method well-approximates the solution derived with the continuous switching. We also evaluate an element of covariance matrix $\braket{Q^{2}(t)}$ in Fig. 3(b) when the system is initially prepared in a coherent state. As $G(t)$ is already well-approximated by $G_{f_{P}}(t)$, the correlation function $\braket{Q^{2}(t)}$ is also approximated by delta switchings when $N$ is large enough.
V Conclusion
We considered the QLE in the Caldeira-Leggett model with a time-dependent coupling and showed that a train of Dirac delta switchings is a powerful tool to deal with the equation. In particular, the delta switchings allow us (i) to solve the QLE exactly even if the dissipation kernel is not time-translation invariant, and (ii) to naturally incorporate the notion of discretized data (i.e., the samples of continuous functions) so that the DTFT can be employed for computing two-point correlation functions.
The solution to the QLE using the delta switchings is expressed as a sum of all possible influences of the memory effect in an environment. This can be intuitively understood by using diagrams in Figs. 1 and 2, which can also help us to construct the solution $Q(t)$ without actually solving the equation.
Our method can also be applied to the QLE in the relativistic settings. In this case, the system is modeled by a harmonic oscillator-type Unruh-DeWitt particle detector [23, 24] coupled to a quantum field in curved spacetime. However, it is crucial to introduce a smearing function (i.e., a detector’s size) to avoid the UV divergences due to the nature of delta-switchings. This is currently under investigation by the authors.
Acknowledgements.
K.G.Y. is supported by Grant-in-Aid for JSPS Fellows Grant Number JP25KJ0048. TT gratefully acknowledges the support of K2-Spring for providing the resources necessary for this research.
Appendix A Markovianity and the covariance matrix
A.1 Transformation of covariance matrices
As described in Sec. III.3, Gaussian states are fully determined by the first moment $\bar{\bm{R}}\equiv\braket{\bm{R}}$ and the covariance matrix $\mathbb{V}$. Any tensor product of Gaussians is Gaussian, and Gaussian operations are those CP maps preserving the Gaussianity. In particular, the full Hamiltonian in the Caldeira-Leggett model generates a Gaussian unitary. If the initial joint state is a product of Gaussian states, $\rho_{\text{tot}}(0)=\rho_{\text{s}}(0)\otimes\rho_{\text{E}}(0)$, then the final reduced state of the system, $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U_{t}\rho_{\text{tot}}(0)U_{t}^{\dagger}]$, after the Gaussian unitary time-evolution $U_{t}\equiv U(t,0)$ remains Gaussian.
Since Gaussian states are characterized by covariance matrices, we can equivalently track the evolution of $\mathbb{V}(t)$. In phase space in which covariance matrices are defined, tensor products $\otimes$ become direct sums $\oplus$, and Gaussian unitaries correspond to symplectic matrices. Therefore, computing the system’s final density matrix $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U_{t}\rho_{\text{tot}}(0)U_{t}^{\dagger}]$ is equivalent to computing the system’s covariance matrix $\mathbb{V}(t)$ by performing a partial trace on the final joint covariance matrix $\mathbb{S}_{t}(\mathbb{V}(0)\oplus\mathbb{V}_{\text{E}}(0))\mathbb{S}_{t}^{\intercal}$, where $\mathbb{S}_{t}$ is the symplectic matrix corresponding to the Gaussian unitary operator $U_{t}$ and $\mathbb{V}_{\text{E}}(0)$ is the environment’s initial covariance matrix.
It is well-known that if the joint state is initially separable Gaussian, then the system’s covariance matrix $\mathbb{V}$ evolves as [20]
$\mathbb{V}(t)$ $=\mathbb{T}_{t}\mathbb{V}(0)\mathbb{T}^{\intercal}_{t}+\mathbb{N}_{t}\,,$ (56)
where $\mathbb{T}_{t}\equiv\mathbb{T}(t,0)$ and $\mathbb{N}_{t}\equiv\mathbb{N}(t,0)$ are $2\times 2$ matrices that transform the system’s covariance matrix at $t=0$ to time $t$. However, if the initial joint state is not separable, no such simple affine form in terms of $\mathbb{V}(0)$ alone exists.
As an example, suppose we apply two Gaussian unitaries $U_{t}\equiv U(t,0)$ and $U_{T-t}\equiv U(T,t)$ to a separable initial joint Gaussian state $\rho_{\text{tot}}(0)$. The system’s final state, $\rho_{\text{s}}(T)=\operatorname{Tr}_{\text{E}}[U_{T-t}U_{t}\rho_{\text{tot}}(0)U^{\dagger}_{t}U^{\dagger}_{T-t}]$, is Gaussian, and the corresponding covariance matrix of the system can be expressed as
$\mathbb{V}(T)$ $=\mathbb{T}_{T}\mathbb{V}(0)\mathbb{T}^{\intercal}_{T}+\mathbb{N}_{T}\,,$
where $\mathbb{T}_{T}$ and $\mathbb{N}_{T}$ are square matrices corresponding to the entire Gaussian time-evolution $U_{T-t}U_{t}(\equiv U_{T})$. Similarly, for the first evolution $U_{t}$ alone,
$\mathbb{V}(t)$ $=\mathbb{T}_{t}\mathbb{V}(0)\mathbb{T}^{\intercal}_{t}+\mathbb{N}_{t}\,.$
Here, $\mathbb{T}_{t}\equiv\mathbb{T}(t,0)$ and $\mathbb{N}_{t}\equiv\mathbb{N}(t,0)$ with $t\in(0,T)$ correspond to the time-evolution under $U_{t}$. However, the following relation generally does not hold:
$\mathbb{V}(T)$ $=\mathbb{T}_{T-t}\mathbb{V}(t)\mathbb{T}_{T-t}^{\intercal}+\mathbb{N}_{T-t}\,,$ (57)
where $\mathbb{T}_{T-t}\equiv\mathbb{T}(T,t)$ and $\mathbb{N}_{T-t}\equiv\mathbb{N}(T,t)$ are meant to transform the covariance matrix $\mathbb{V}(t)$ to $\mathbb{V}(T)$. This is because the system entangles with the environment after the first Gaussian unitary $U_{t}$, thereby the actual expression that relates $\mathbb{V}(t)$ to $\mathbb{V}(T)$ is more complicated. This fact is relevant in our Dirac-delta method as the entire Gaussian unitary operator can be decomposed into a product $U_{T}=U_{T-t_{N-1}}U_{t_{N-1}-t_{N-2}}\ldots U_{t_{1}}$, where each unitary operator is generated by a delta-coupling.
A.2 Transformation of covariance matrices under the Markovian dynamics
Although Eq. (57) does not hold in general, we will show that it does for Markovian dynamics. To this end, we introduce the concepts of universal dynamical maps (UDMs) and Markovian dynamical maps.
The quantum Markovianity of a quantum dynamical map is often characterized by the completely positive trace-preserving (CPTP) property, along with the phenomenologically motivated semigroup property, which ensures an irreversible dynamical process. These properties are equivalent to the dynamical map satisfying the well-known GKSL master equation [2], [^6] The CPTP and semigroup properties are equivalent to the generalized GKSL master equation, where the Hamiltonian, coupling, and Kraus operators can depend on time. forming the foundation of Markovian quantum dynamics.
In this work, we define Markovian dynamics as those that can be expressed as a CPTP semigroup map when the initial condition is a Gaussian state. However, we emphasize that this is not the only definition of Markovianity, as various alternative characterizations exist [25].
Let us begin by defining Markovianity [26]. Consider a quantum channel (i.e., a CPTP map) that maps a system’s arbitrary quantum state from $t=0$ to time $t$, $\mathcal{E}_{t,0}:\rho_{\text{s}}(0)\mapsto\rho_{\text{s}}(t)$. Quantum channels can be utilized even when a system is initially entangled with an environment. A universal dynamical map (UDM) is a special class of quantum channels, where the system’s input state is assumed to be uncorrelated with the environment. For example, a map induced by $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U_{t}\rho_{\text{tot}}(0)U_{t}^{\dagger}]$ is a UDM if the system is initially uncorrelated with the environment.
A Markovian dynamics can be characterized in terms of the divisibility of a UDM. Suppose a map $\mathcal{E}_{t,0}$ is a UDM and it can be decomposed into two maps:
$\mathcal{E}_{t,0}=\mathcal{E}_{t,s}\circ\mathcal{E}_{s,0}\,,\quad\forall t\geq s\geq 0\,.$ (58)
In general, however, it is not necessarily true that both $\mathcal{E}_{t,s}$ and $\mathcal{E}_{s,0}$ are UDMs. For example, $\mathcal{E}_{s,0}$ can be a UDM, but $\mathcal{E}_{t,s}$ is generally not, as the first UDM $\mathcal{E}_{s,0}$ entangles the system and the environment. A UDM $\mathcal{E}_{t,0}$ is called a Markovian dynamical map if both maps $\mathcal{E}_{t,s}$ and $\mathcal{E}_{s,0}$ are UDMs. This is also known as the one-parameter semigroup property, and it can be shown that it is equivalent to the generalized GKSL equation [26].
Next, we consider how the above Markovianity condition applies to the formulation using covariance matrices in Gaussian states. In particular, we show that if the time-evolution is described by a Markovian dynamical map, then the matrices $\mathbb{T}$ and $\mathbb{N}$ satisfy
$\mathbb{T}_{t}=\mathbb{T}_{t-s}\mathbb{T}_{s}\,,\quad\mathbb{N}_{t}=\mathbb{T}_{t-s}\mathbb{N}_{s}\mathbb{T}^{\intercal}_{t-s}+\mathbb{N}_{t-s}\,,$ (59)
where $\mathbb{T}_{t-s}\equiv\mathbb{T}(t,s)$ and $\mathbb{N}_{t-s}\equiv\mathbb{N}(t,s)$ are the matrices associated with a UDM $\mathcal{E}_{t,s}$.
To show this, consider the system’s initial Gaussian state $\rho_{\text{s}}(0)\equiv\rho_{\text{s}}[\bar{\bm{R}}(0),\mathbb{V}(0)]$. Assuming the joint initial state is a Gaussian product state and that the entire system evolves under a Gaussian unitary, the system’s state at time $t$ is expressed as
$\mathcal{E}_{t,0}[\rho_{\text{s}}[\bar{\bm{R}}(0),\mathbb{V}(0)]]=\rho_{\text{s}}[\mathbb{T}_{t}\bar{\bm{R}}(0),\mathbb{T}_{t}\mathbb{V}(0)\mathbb{T}_{t}^{\intercal}+\mathbb{N}_{t}]\,,$ (60)
where $\mathcal{E}_{t,0}$ is a UDM and we used the relation (33). Suppose the UDM $\mathcal{E}_{t,0}$ is a Markovian dynamical map. Then, we have the following relations:
$\rho_{\text{s}}[\bar{\bm{R}}(s),\mathbb{V}(s)]$ $=\mathcal{E}_{s,0}[\rho_{\text{s}}[\bar{\bm{R}}(0),\mathbb{V}(0)]]$
$=\rho_{\text{s}}[\mathbb{T}_{s}\bar{\bm{R}}(0),\mathbb{T}_{s}\mathbb{V}(0)\mathbb{T}_{s}^{\intercal}+\mathbb{N}_{s}]\,,$
$\mathcal{E}_{t,s}[\rho_{\text{s}}[\bar{\bm{R}}(s),\mathbb{V}(s)]]$ $=\rho_{\text{s}}[\mathbb{T}_{t-s}\bar{\bm{R}}(s),\mathbb{T}_{t-s}\mathbb{V}(s)\mathbb{T}_{t-s}^{\intercal}+\mathbb{N}_{t-s}]\,.$
Thus, the semigroup property $\mathcal{E}_{t,0}=\mathcal{E}_{t,s}\circ\mathcal{E}_{s,0}$ leads to (59), which completes our proof.
Appendix B Proof of Markovianity in the Limit $\Lambda T/N\gg 1$ and $\gamma/\Omega\ll 1$
In this section, we demonstrate that the two approximations, $\Lambda T/N\gg 1$ and $\gamma/\Omega\ll 1$, imply the Markovianity condition (59). We refer to these as the Markov and Born approximations, respectively.
The explicit forms of the transformation matrices $\mathbb{T}_{t}$ and $\mathbb{N}_{t}$ are given by
$\mathbb{T}_{t}$ $=\begin{bmatrix}G_{f_{Q}}(t)&G_{f_{P}}(t)\\ \dot{G}_{f_{Q}}(t)&\dot{G}_{f_{P}}(t)\end{bmatrix},$ (61)
$\mathbb{N}_{t}$ $=\begin{bmatrix}N_{QQ}(t)&N_{QP}(t)\\ N_{PQ}(t)&N_{PP}(t)\end{bmatrix},$ (62)
where the functions $G_{f_{Q}}(t)$ and $G_{f_{P}}(t)$ are defined in Eq. (24), and the matrix components $N_{QQ}(t)$, $N_{QP}(t)$, $N_{PQ}(t)$, and $N_{PP}(t)$ are specified in Eq. (38).
Next, we recall that the quantity $G_{f}(t)$ can be expressed as a combination of the function $f$ and a matrix polynomial in $K$, as defined in Eq. (18). Since the matrix $K$ is of order $\mathcal{O}(\gamma/\Omega)$, the first-order Born approximation of $G_{f}(t)$ is given by
$G_{f}(t)=f(t)+\sum_{l=1}^{N-1}K_{t,t_{l}}f(t_{l})+\mathcal{O}\left((\gamma/\Omega)^{2}\right),$ (63)
for $f=f_{Q}$ and $f=f_{P}$. Note that the term with $l=N$ does not contribute to the sum above, since $K_{t,t_{N}}=0$ when $t=t_{N}$.
The explicit form of $K_{t,t_{l}}$ is given by
$K_{t,t_{l}}=$ $\sum_{k=1}^{N}\gamma\left(\frac{\Lambda T}{N}\right)^{2}e^{-\frac{\Lambda T}{N}|k-l|}\Theta(t_{k}-t_{l})$
×$\times\frac{\sin\left[\Omega(t-t_{k})\right]}{\Omega}\Theta(t-t_{k}).$ (64)
Under the Markov approximation, $\Lambda T/N\gg 1$, the leading-order term in the above summation is $k=l+1$:
$K_{t,t_{l}}=$ $\gamma\left(\frac{\Lambda T}{N}\right)^{2}e^{-\Lambda\frac{T}{N}}\frac{\sin\left[\Omega(t-t_{l+1})\right]}{\Omega}$
$+\mathcal{O}\left(\left(\frac{\Lambda T}{N}\right)^{2}\left(e^{-\frac{\Lambda T}{N}}\right)^{2}\right).$ (65)
Thus, under the combined Born-Markov approximation, the function $G_{f}(t)$ becomes
$G_{f}(t)\simeq$ $f(t)+\sum_{l=1}^{N-1}\frac{T^{2}}{N^{2}}\Gamma\left(\frac{T}{N}\right)\frac{\sin\left[\Omega(t-t_{l+1})\right]}{\Omega}f(t_{l})\,,$ (66)
where $\Gamma(t)$ denotes the Lorentz-Drude spectral density defined in Eq. (46).
This expression indicates that, under the Born-Markov approximation, the leading-order contribution arises from propagation over a single time step interval, i.e., from $t_{l}$ to $t_{l+1}$. Propagations over longer intervals, such as from $t_{l}$ to $t_{l+n}$ for $n=2,3,\dots$, are exponentially suppressed by terms of the form $(\Lambda T/N)^{2}\exp(-n\Lambda T/N)$, where $n$ is the number of time steps. Therefore, retaining only the one-time-step propagation diagram, as illustrated in Fig. 2, corresponds to the Markovian dynamics.
We now proceed to verify the semigroup property $\mathbb{T}_{t}=\mathbb{T}_{t-s}\mathbb{T}_{s}$. Since $t>s>0$, without loss of generality, let us denote $s=t_{j}$ for some $N>j>0$. Then, we aim to show the matrix identity:
$\begin{bmatrix}G_{f_{Q}}(t)&G_{f_{P}}(t)\\ \dot{G}_{f_{Q}}(t)&\dot{G}_{f_{P}}(t)\end{bmatrix}=\begin{bmatrix}G_{f_{Q}}(t-t_{j})G_{f_{Q}}(t_{j})+G_{f_{P}}(t-t_{j})\dot{G}_{f_{Q}}(t_{j})&G_{f_{Q}}(t-t_{j})G_{f_{P}}(t_{j})+G_{f_{P}}(t-t_{j})\dot{G}_{f_{P}}(t_{j})\\ \dot{G}_{f_{Q}}(t-t_{j})G_{f_{Q}}(t_{j})+\dot{G}_{f_{P}}(t-t_{j})\dot{G}_{f_{Q}}(t_{j})&\dot{G}_{f_{Q}}(t-t_{j})G_{f_{P}}(t_{j})+\dot{G}_{f_{P}}(t-t_{j})\dot{G}_{f_{P}}(t_{j})\end{bmatrix}.$ (67)
The right-hand side can be simplified using trigonometric identities, such as
$\cos(\Omega t)=\cos[\Omega(t-t_{j})]\cos(\Omega t_{j})-\sin[\Omega(t-t_{j})]\sin(\Omega t_{j}),$ (68)
$\sin[\Omega(t-t_{l+1})]=\cos[\Omega(t-t_{j})]\sin[\Omega(t_{j}-t_{l+1})]+\sin[\Omega(t-t_{j})]\cos[\Omega(t_{j}-t_{l+1})].$ (69)
By applying these identities, one can verify the equality of the $(1,1)$ and $(1,2)$ entries of the matrix in Eq. (67). The equalities of the $(2,1)$ and $(2,2)$ components then follow by taking the $t$-derivatives of the corresponding $(1,1)$ and $(1,2)$ entries.
Thus, the equality in Eq. (67) holds to first order in $\gamma/\Omega$.
Similarly, the noise matrix satisfies the composition law $\mathbb{N}_{t}=\mathbb{T}_{t-s}\mathbb{N}_{s}\mathbb{T}_{t-s}^{\intercal}+\mathbb{N}_{t-s}$, which can be shown by employing the same trigonometric identities as in the proof above.